# DoubleDiffusion: Combining Heat Diffusion with Denoising Diffusion for Generative Learning on 3D Meshes

### v3

@@ Abstract
This paper proposes DoubleDiffusion, a novel framework that combines heat dissipation diffusion and denoising diffusion for direct generative learning on 3D mesh surfaces. Our approach addresses the challenges of generating continuous signal distributions residing on a curve manifold surface. Unlike previous methods that rely on unrolling 3D meshes into 2D or adopting field representations, DoubleDiffusion leverages the Laplacian-Beltrami operator to process features respecting the mesh structure. This combination enables effective geometry-aware signal diffusion across the underlying geometry. As shown in Fig. 1, we demonstrate that DoubleDiffusion has the ability to generate RGB signal distributions on complex 3D mesh surfaces and achieves per-category shape-conditioned texture generation across different shape geometry. Our work contributes a new direction in diffusion-based generative modeling on 3D surfaces, with potential applications in the field of 3D asset generation. https://github.com/Wxyxixixi/DoubleDiffusion_3D_Mesh
Figure 1: We present DoubleDiffusion, a framework for generative learning on 3D manifold mesh. We use (a) the heat dissipation diffusion as the geometric feature extractor to efficiently aggregate vertice features over the spatial domain on mesh surfaces. We employ a (b) temporal denoising diffusion process to capture the distribution of the signal on mesh surfaces. Our model can effectively learn the RGB signal distribution defined on a specific manifold (e.g., (c) sinewave and (d) bunny), to generate different (c) handwritten digits [11] or (d) face looks [18]. Moreover, Our model can generate the texture on different manifold shape given different manifold surfaces from the (e) ShapeNetCore [7].
@@ 1 Introduction
Significant advancements have been made recently in generative learning using the diffusion probabilistic models [14, 33], particularly in domains involving images [29, 26], videos [15, 6], and 3D such as points clouds [23], triplanes [32], Nerf [21], mesh texture generation based on grid [22], etc. However, the extension of generative learning to 3D object surfaces, typically represented as curved surfaces or manifolds, remains under-explored due to the lack of effective learning strategies on these warped manifold surfaces. Several existing methods [8, 42] attempt to learn the texture distribution on 3D object manifold surfaces through unwarped 2D texture maps. While these approaches enable surface representation learning in 2D format and can leverage 2D image models to infer texture samples, they often suffer from the inconsistency along edges and fail to fully capture the intricate geometry of the 3D surface. Additionally, unwrapping a 3D surface onto a 2D texture map is ill-posed without a certain solution, thus different unwrapping techniques can significantly impact the final generative results.
While some methods used the point cloud as the manifold surface representation, the neighborhood relationship and the distance are computed through K-nearest neighbors (KNN), the distance are not ideally calculated along the manifold surface [27, 28]. Meanwhile, this approach lacks the precision needed for mesh-based structures, where each point’s neighboring relationships are defined by the mesh’s inherent geometry. Since the mesh is naturally a structured manifold, with the geodesic distance represented via the vertices connectivity, mesh present an ideal medium for diffusion-based generation directly on the surface.
In this work, we present a probabilistic diffusion framework on manifold, addressing the challenge of generative learning directly on the mesh surfaces. Our method captures the local geometric characteristics of 3D shapes without needing to flatten the surface to a 2D medium, ensuring consistency of the content generation across the manifold. Bronstein [5] points out that, Laplacian Beltrami (LB) operator can capture the differential geometric property on the manifold defining neighboring and distance relationship when the information interchanging among the points. To model the diffusion on the manifold mesh surface, we adapt from the mesh processing theory [43] and use the Laplacian Beltrami operator to capture the local geometric of the manifold mesh. Since mesh is a structured manifold, where the geometry of the manifold are explicitly entailed in the mesh topology, the geometric information of the manifold mesh, such as curvatures, geodesic distances of the mesh can be captured and process through the discrete Laplacian-Beltrami Operators [34, 43], opening up a great opportunity to directly learn a scalar distribution defined on the manifold domain.
Specifically, we employ the LB operator as a core feature processing element to approximate heat dissipation diffusion. This diffusion process models the principles of heat transfer on continuous surfaces, where heat spreads over time based on the geometry of the surface, allowing information to propagate across points in a structured manner [13, 9, 31]. On a mesh manifold, the LB operator can approximate this process, effectively simulating heat diffusion and enabling smooth information transfer across vertices. In this paper, we propose to integrate the heat dissipation diffusion with the diffusion probabilistic model [14] as the generative learning framework on the manifold. To our knowledge, we are the first work to combine heat dissipation diffusion with a denoising diffusion probabilistic model (DDPM) for direct learning signal distributions defined on the manifold mesh. We term our framework DoubleDiffusion, as it leverages heat dissipation diffusion to capture surface-specific features and employs the diffusion probabilistic framework for generative learning directly on the manifold. In summary, our contributions are:
  1. 1.
@@
  2. 2.
    Our framework achieves a significant 312.82% improvement in coverage compared with the previous SoTA MDF [12]. Meanwhile, it can process large meshes of size e.g., $\sim$100k vertices in a single pass,$\sim$20x more than MDF. It is also 8.1x faster to generate a sample on the manifold mesh.
  3. 3.
@@ 3.1 Heat Diffusion
Heat diffusion has long been used in describing how an inital scalar heat distribution $\bm{h}$ diffuses or “spread out” over time ‘$s$’. In the differential geometry, it can describe how the signal diffuse over the continuous surfaces. Mathematically, this process is defined as a partial differential equation (PDE):
@@
where $\bm{h}(s)$ represents the feature energy (or heat) over the manifold, $s$ represent the distance of how far the energy is transmitted, and $\Delta$ is the Laplacian Beltrami operator characterizes how the tempreature at a point dicerges from its surrounding region.
@@ 3.2 Mesh Laplacian
In the context of 3D mesh, that is, a discrete representation of surface manifolds composed of a finite number of vertices $\mathbf{V}$ and triangular faces $\mathbf{F}\subset\mathbf{V}^{3}$, we dicretize the continuous heat equation to operate on the mesh structure. And we use the cotangent-weighted Mesh Laplacian $L^{MC}$ proposed in [24] as a common practice.
$\mathbf{L}=M^{-1}\mathbf{C}.$ (2)
The cotangent weight matrix $C$ is calculated with respect to the angles adjacent to the shared edges, and the mass matrix is related to the area of the adjacent faces. These reflect the differential relationships and local geometry among vertices, acting similarly to an integration operation over the geodesic neighborhood around each vertex on manifold surfaces [4].
The cotangent weights $C$ are calculated as follows:
$C_{ij}=\sum_{j\in N(i)}\frac{\cot\alpha_{ij}+\cot\beta_{ji}}{2},$ (3)
where $C_{ij}\in\mathbb{R}$ represents the cotangent weight along the edge $(i,j)$, and $\alpha_{ij}$ and $\beta_{ji}$ are the angles opposite to edge $\bm{e}_{ij}$. This formulation considers both edge lengths and adjacent triangle areas (Fig.2), which are loosely related to the Gaussian curvature at the central vertex $v_{i}$, as stated in [24].
The cotangent is calculated by:
$\cot\alpha_{ij}$ $=\frac{\bm{e}_{ki}\cdot\bm{e}_{jk}}{|\bm{n}_{k}|},$ (4)
$\bm{e}_{ki}=\bm{v}_{i}-\bm{v}_{k},\bm{e}_{jk}$ $=\bm{v}_{j}-\bm{v}_{k},n_{k}=\bm{e}_{i}\times\bm{e}_{j},$
where $\bm{e}_{i}$ and $\bm{e}_{j}$ are the edge vectors.
When angles exceed $\frac{\pi}{2}$ , cotangent values become negative. To ensure a positive semi-definite Laplacian matrix, we follow common practice by applying a mass matrix $M\in\mathbb{R}^{n\times n}$, $M$ is a mass matrix given by: $M_{ii}=\frac{1}{3}\sum_{i\in t}\text{Area}(t),$ which assigns one-third of the area of each incident triangle to vertex $\bm{v}_{i}$, thus accounting for the vertex’s share of the surrounding area. Having the mass matrix into the cotan-weight Laplacian matrix makes it an FEM Laplacian operator and ensures that its eigenvalue decomposition is robust to mesh parameterization, as discussed in [43].
@@ 4 Method
4.1 Overview
                 (a)           (b)                      (c)                            (d)
Figure 3: Framework Overview. Our (a) DoubleDiffusion Network contains consecutive (b) DoubleDiffusion Blocks (DDB). Each DDB consists of the (c) Heat Diffusion module aggregating vertice features over the spatial domain on mesh surfaces and (d) the Timestep-Aware MLP that injects timestep embedding in a residual manner.
Our method begins by computing the cotangent-weighted Laplacian $\mathbf{L}$ from the manifold mesh structure (Sec.3.2), which serves as the foundation for diffusion on the manifold. Using this Laplacian, we perform eigen decomposition on the Laplacian-Beltrami operator, obtaining the smallest $k$ eigenvalues and eigenvectors.These are then used to approximate a heat diffusion kernel $\mathbf{H}(s)$across the manifold (Sec.4.2).
This heat dissipation diffusion acts as a geometric feature extractor on the manifold, where $s$ in $\mathbf{H}(s)$ represents the distance that a feature diffuses from vertex $\bm{v}_{i}$ across the manifold mesh $\mathcal{M}$. Analogous to a convolutional kernel on an image grid, this heat dissipation diffusion aggregates features on the scalar function $\bm{f}$ defined on the vertices of the manifold mesh as $\mathbf{L}\bm{f}$. Here, $\bm{f}=(\bm{f}^{v_{0}},\bm{f}^{v_{1}},\cdots,\bm{f}^{v_{n}})$, where $\bm{f}^{v_{1}}$ refers to the scalar vector on vertex $\bm{v}_{i}$, and $\{\bm{v}_{0},\bm{v_{1}},\cdots,\bm{v_{n}}\}\in\mathcal{M}$.
Our diffusion probabilistic model generates samples by starting with a noise distribution $\bm{f}_{T}^{v_{i}}\sim\mathcal{N}(0,1)$, the initial noise is independently sampled at each vertex. The DPM is trained to iteratively predict the reverse transitions $p(\bm{f}_{T}^{v_{i}}|\bm{f}_{T-1}^{v_{i}})$ progressively denoising through each step.
It’s worth noting that, we define two independent parameters: the heat dissipation diffusion distance $s$ and the denoising diffusion time step $t$ in the DoubleDiffusion framework. The diffusion distance $s$ is a learnable parameter within the heat dissipation network, allowing the model to adaptively capture geometric features on the manifold. In contrast, the denoising diffusion time step $t$ follows a predefined schedule, iteratively progressing from $T$ to 00 as part of a recursive Markov chain learning process.
4.2 Laplace Approximation of Heat Diffusion
Solving the heat equation presents an eigenvalue problem, where we can use the mesh Laplacian in Eq. 2 to approximate the heat diffusion. This approach also brings effect to fast compute to the heat diffusion Eq. 1 on large meshes. Meanwhile, Laplacian operators has been recognized as a spectral filer [37, 20]. Choosing the eigenvector of the mesh Laplacian $\mathbf{L}$ with $k$ smallest eigenvalues, Laplacian operator selectively attenuates high-frequency components (sharp changes or noise) while preserving low-frequency components (smooth, global variations). Thus, the Laplacian-approximated heat diffusion functions by dissipating high-frequency variations more quickly than low-frequency components. This behavior effectively smooths the signal, making the heat diffusion process an ideal tool for filtering.
We use the eigen-decomposition of the Laplacian operator in Eq. 2, yielding:
@@ 4.2 Laplace Approximation of Heat Diffusion
where $\Phi$ contains the eigenvectors of $L$, and $\Lambda$ is a diagonal matrix with eigenvalues $\lambda_{i}$. The heat diffusion is then approximated as:
@@
In this spectral representation, the heat diffusion kernel $e^{-s\Lambda}$ has diagonal elements $e^{-s\lambda_{i}}$ that control the rate of decay for each frequency component.
Furthermore, we followed the method introduce in [31], to apply the heat kernal with respect to the feature channel:
@@
where the $\Phi$ only contains the $k$ smallest eigenvector of the $L^{MC}$ and the eigenvalues are in small to large order $\lambda_{0}<\lambda_{1}<\cdots<\lambda_{k}$. The symbol $\odot$ represents the element-wise multiplication. In this process, the $k$-eigenvectors first project the feature signals into the spectral domain, where the high-frequency components are suppressed by the frequency-modulated heat kernel, while low-frequency components are enhanced. This approach achieves channel-wise feature aggregation across vertices in the spectral domain, effectively gathering features within the heat diffusion range defined by the learnable $s$.
This connection between the heat diffusion equation and low-pass filtering allows us to effectively aggregate features across the manifold while preserving geometric consistency.
Denoising Process with Heat Diffusion. The reverse process $p_{\theta}(\bm{f}_{t-1}|\bm{f}_{t})$ removes noise iteratively. We parameterize this by predicting the noise $\bm{\epsilon}_{\theta}(\bm{f}_{t},t)$ using the heat diffusion kernel $\mathbf{H}(s)$, where $s$ is the heat diffusion distance:
@@
where $\mu_{\theta}(\bm{f}_{t},t)=\frac{1}{\sqrt{1-\beta_{t}}}\left(\bm{f}_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\bm{\epsilon}_{\theta}(\bm{f}_{t},t)\right)$.
Following previous work [30], we incorporate the time step $t$ as a time embedding within each consecutive block using a per-vertex MLP. Given that heat diffusion is conducted on a per-channel basis, employing group normalization [40] directly disrupts the preservation of spatial information, resulting in non-convergence of the model. Consequently, we empirically adopt layer normalization [2] to enhance the training process.

### v5

# DoubleDiffusion: Combining Heat Diffusion with Denoising Diffusion for Texture Generation on 3D Meshes

@@ Abstract
This paper addresses the problem of generating textures for 3D mesh assets. Existing approaches often rely on image diffusion models to generate multi-view image observations, which are then transformed onto the mesh surface to produce a single texture. However, due to the gap between multi-view images and 3D space, such process is susceptible to a range of issues such as geometric inconsistencies, visibility occlusion, and baking artifacts. To overcome this problem, we propose a novel approach that directly generates texture on 3D meshes. Our approach leverages heat dissipation diffusion, which serves as an efficient operator that propagates features on the geometric surface of a mesh, while remaining insensitive to the specific layout of the wireframe. By integrating this technique into a generative diffusion pipeline, we significantly improve the efficiency of texture generation compared to existing texture generation methods. We term our approach DoubleDiffusion, as it combines heat dissipation diffusion with denoising diffusion to enable native generative learning on 3D mesh surfaces.
https://github.com/Wxyxixixi/DoubleDiffusion_3D_Mesh
@@ 1 Introduction
Texture generation task takes in an uncolored 3D object and output the texture color on the 3D object. The input object can be any of the 3D representations, such as Nerf, Gaussian Splatting, Point Clouds, Mesh, etc. While many methods target at Nerf [16, 19] and Gaussian Splatting [27] based texture generation, fewer methods have explored the texture generation on 3D mesh. Since meshes are widely used as a 3D representation in graphic design, generating texture directly on a given mesh has become an essential approach.
Due to the lack of a unified 3D foundation model for meshes, some prior works [5, 35] unwrap the surface into 2D texture maps and harness powerful image priors such as Stable Diffusion [20]. Although UV maps lie on a 2D plane, they are having geometric distortion and the texture cannot be generated directly via a 2D diffusion model. To bridge this gap, some methods takes the multi-view texture generation method as remedy, by using the 2D view images as the guidance for the objects’ renders with the generated texture [5, 3, 29, 24]. Despite leveraging strong 2D priors, these methods require considerable computational resources to maintain multi-view consistency and often suffer from edge inconsistencies. Furthermore, these methods are not native mesh approach for texture generation.
Figure 1: Overview of the DoubleDiffusion.
In our method, we frame texture generation on a 3D mesh as a signal generation problem on a curved surface. Unlike image generation on flat grids, the curvature of the geometric mesh plays a crucial role, making it impractical to directly apply a standard 2D convolutional operator. Although modern diffusion models have made significant progress in image domains – thanks to convolutional or sequential networks operating on regular grids – it remains challenging to transfer these convolutional kernels or auto-regressive modules directly to meshes. While some methods [7, 8] address this by applying positional encoding to sampled points to impose an order for signal learning, our approach instead leverages the native vertex-edge data structure.
Generating signals directly on the mesh, however, poses its own challenges. Mesh data structures are irregular, and each vertex’s neighborhood relationships lack a simple pattern, demanding a unified operator that respects the underlying geometry. We treat texture generation on the mesh as learning a color distribution anchored in the mesh’s geometry. Since the mesh is naturally a structured manifold where geodesic distances are defined by vertex connectivity, manifold operators are well-suited to capturing signal distributions on its surface. In mesh processing, the Laplacian-Beltrami (LB) operator [36] serves as a second-order differential operator that measures how a scalar function varies with respect to the surface geometry. Its discrete form can process signal variation on the mesh, and Sorkine [25] further shows that applying the discrete LB operator to the vertices corresponds to their mean curvature. Using the cotangent Laplacian, which encodes differential coordinates, effectively represents the local relationships between a vertex and its neighbors.
By viewing the texture signals on the manifold as a form of “heat energy,” we can apply the heat diffusion kernel to capture how signals evolve with respect to geometric neighbors. This concept parallels using convolutional or transformer layers in image-based diffusion models, except that our “noise predictor” in the denoising diffusion pipeline is designed for mesh geometry. Specifically, we embed a heat-diffusion-based network, DiffusionNet [23], into a denoising diffusion framework to model RGB signal distributions directly on the mesh surface.
Albeit this straightforward design, direct operation on mesh proves highly efficient at learning RGB distributions while maintaining fast computation. We evaluate its performance on two types of generation tasks. First, we showcase how our method handles complex signal distributions on a single curved manifold, illustrating how geometry-aware diffusion enhances texture generation. Second, we test on the ShapeNet Core Dataset, which contains 50 categories, each with thousands of shapes, to assess the method’s capacity to learn textures across varying geometry and category-specific appearances. Notably, our pipeline processes batches of different shapes, learning the joint distribution of geometry and category-constrained texture.
In summary, our key contributions are:
@@
  • We compare with the current SOTA manifold based generation method MDF [7] and achieves a significant 312.82% improvement in coverage. Meanwhile, it can process large meshes of size e.g., $\sim$100k vertices in a single pass. It is also 8.1x faster to generate a sample on the manifold mesh.
@@ 3.1 Heat Diffusion
Heat diffusion has long been used in describing how an initial scalar heat distribution $\bm{h}$ diffuses or “spread out” over time ‘$s$’. To distinguish the heat-diffusion time ‘$s$’ from the denoising diffusion timestep $t$, we use $s$ to represent the duration of heat diffusion. In differential geometry, heat diffusion on a mesh surface follows:
@@
where $\bm{h}(s)$ denotes the heat or feature energy across the surface, and $\Delta$ is the Laplacian Beltrami operator that characterizes how temperature at each point diverges from its surroundings. The heat energy spreading speed depends on the design of the Laplacian matrix and the heat function. In our method, we used the DiffusionNet [23], which supports regions of greater curvature or density experience more extensive feature-energy exchange during the diffusion process.
@@ 3.2 Mesh Laplacian
In our method, we adopt the cotangent-weight Laplacian matrix [17] as the Laplacian Beltrami Operator in Eq. 1. A typical mesh consists of vertices, edges, and faces. In graph learning, the graph Laplacian has a long history of usage, focusing primarily on node connectivity. By contrast, the cotangent-weight Laplacian matrix has been widely used as a geometric mesh Laplacian, where the edge weights depend not only on edge lengths but also on the areas enclosed by triangular faces. Consequently, it is loosely related to the Gaussian curvature at the central vertex $v_{i}$ [17].
Applying cotangent-weight Laplacian Beltrami Operator gives:
$[\mathbf{L}\bm{h}]_{i}=\frac{1}{\Omega_{i}}\sum_{j\in N(i)}\frac{1}{2}(\cot\alpha_{i}j+\cot\beta_{i}j)(h_{i}-h_{j})$ (2)
where $\mathbf{L}$ represent the Laplacian Beltrami Operator, the $\Delta$ in Eq. 1. $\bm{h}_{i}$ represent the heat energy that being received for the vertices i. $\Omega_{i}$ is the total area of the triangle that contains vertices i, $N(i)$ are the vertices that is neighbour to the vertices i, and $\alpha_{ij}$, $\beta_{ij}$ represents the angles that opposite to the edge $e_{ij}$.
@@ 4 Method
4.1 Overview of DoubleDiffusion
                                    (a)                                             (b)                                                                                           (c)
Figure 3: Method Overview. Our (a) DoubleDiffusion Network contains consecutive (b) DoubleDiffusion Blocks (DDB). Each (c) DDB consists of the Heat Diffusion module aggregating vertice features over the spatial domain on mesh surfaces and the Timestep-Aware MLP that injects timestep embedding in a residual manner.
The DoubleDiffusion idea is simple and straightforward. First of all, we have an uncolored mesh $\mathcal{M}=(\mathbf{V},\mathbf{F})$ with $n$ number of vertices $\{\bm{v}_{0},\bm{v_{1}},\cdots,\bm{v_{n}}\}\in\mathbf{V}$. And we describe the vertices color or the vertices features as the scalar function $\bm{f}$ resides on the manifold mesh, $\bm{f}=(\bm{f}^{v_{0}},\bm{f}^{v_{1}},\cdots,\bm{f}^{v_{n}})$, where $\bm{f}^{v_{1}}$ refers to the scalar vector on vertex $\bm{v}_{i}$. Here, the $\bm{f}$ is having the similar meaning to the $\bm{h}$ in Eq. 1, in our task, $\bm{f}$ are initially sampled from the Gaussian Noise $\bm{f}_{T}\sim\mathcal{N}(0,1)$. These noisy vertex color will go through the DoubleDiffusion Blocks (DDB) iteratively to be denoised and generate the texture color as shown in Fig. 3a and Fig. 3b.
Each of the DoubleDiffusion Blocks contains a Heat Diffusion and a Timestep-Aware per Vertex MLP as shown in Fig 3. The design of the DDB adapts from DiffusionNet [23], which has been demonstrated as a powerful surface processing network recently [9].
4.2 Heat Diffusion Module
The Heat Diffusion Module contains two main parts. The first part is a geometric-aware low pass filter, which help to spread out the feature to the vertices resides on the geometric mesh. The second part is a spacial gradient calculation, which accelerate the heat diffusion on the surface with respect to the surface gradients.
4.2.1 Geometric-aware Low Pass Filter
The geometric-aware low pass filter as shown in Fig 4 (green) is a Laplacian approximation of the heat diffusion. Solving the heat equation presents an eigenvalue problem, where we can use the eigenvalue decomposition to approximate the heat diffusion. To apply the heat diffusion, we need heat kernel, which is defined as
$h(\lambda_{i})=e^{-s\lambda_{i}}$ (3)
where $s\geq 0$ is a scaling hyper-parameter. The heat kernel can define the convolutional kernel on the surface which is $e^{-sL}$. To compute the heat kernel, we can use the eigen-decomposition to approximate this heat kernel, yielding:
@@ 4.2.1 Geometric-aware Low Pass Filter
where $\Phi$ contains the eigenvectors of $L$, and $\Lambda$ is a diagonal matrix with eigenvalues $\lambda_{i}$.
The heat diffusion on the vertex color is then approximated as:
@@
In this spectral representation, the heat diffusion kernel $e^{-s\Lambda}$ has diagonal elements $e^{-s\lambda_{i}}$ that control the rate of decay for each frequency component. This approximation also brings effect to fast compute to the heat diffusion Eq. 1 on large meshes. Meanwhile, Laplacian operators has been recognized as a spectral filer [30, 15]. Choosing the eigenvector of the mesh Laplacian $\mathbf{L}$ with $k$ smallest eigenvalues, Laplacian operator selectively attenuates high-frequency components (sharp changes or noise) while preserving low-frequency components (smooth, global variations). Thus, the Laplacian-approximated heat diffusion functions by dissipating high-frequency variations more quickly than low-frequency components. This behavior effectively smooths the signal, by discounting high-frequency basic filters via heat kernels.
Furthermore, we followed the method introduce in [23], to apply the heat kernel with respect to the feature channel:
@@
where the $\Phi$ only contains the $k$ smallest eigenvector of the $L^{MC}$ and the eigenvalues are in small to large order $\lambda_{0}<\lambda_{1}<\cdots<\lambda_{k}$. The symbol $\odot$ represents the element-wise multiplication. In this process, the $k$-eigenvectors first project the feature signals into the spectral domain, where the high-frequency components are suppressed by the frequency-modulated heat kernel, while low-frequency components are enhanced. This approach achieves channel-wise feature aggregation across vertices in the spectral domain, effectively gathering features within the heat diffusion range defined by the learnable $s$ (Fig. 4).
Figure 4: Overview of the DoubleDiffusion Block, containing (a) Geometric-aware Low Pass Filter; (b) Spacial gradient Feature and (c) Timestep-aware MLP.
4.2.2 Spatial Gradient Calculations
Following the DiffusionNet, the diffused feature for each vertex will then be multiplied by the gradient operators. The structure of this module is shown in Fig. 4. The calculation of the gradient operators gradX and gradY matrix is calculated with respect to the vertex normals, and the details could be found in  [23].
4.3 Denoising Process with Heat Diffusion.
The DoubleDiffusion generate samples by starting with a noise distribution $\bm{f}_{T}^{v_{i}}\sim\mathcal{N}(0,1)$, the initial noise is independently sampled at each vertex. And the linear noise scheduler [10] is used to prepare the noisy data. Then the DoubleDiffusion is served as the noise predictor $\bm{\epsilon}_{\theta}(\bm{f}_{t},t)$ and being trained to progressively removes noise through each step $p_{\theta}(\bm{f}_{t-1}|\bm{f}_{t})$.
@@ 4.2.2 Spatial Gradient Calculations
$\mu_{\theta}(\bm{f}_{t},t)=\frac{1}{\sqrt{1-\beta_{t}}}\left(\bm{f}_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\bm{\epsilon}_{\theta}(\bm{f}_{t},t)\right)$ (8)
Following previous work [22], we incorporate the time step $t$ as a time embedding within each consecutive block using a per-vertex MLP. Given that heat diffusion is conducted on a per-channel basis, employing group normalization [33] directly disrupts the preservation of spatial information, resulting in non-convergence of the model. Consequently, we empirically adopt layer normalization [2] to enhance the training process as shown in Fig 4(c).